OpenStudy (tw101):

Write the equation of the parabola with a focus (1,2) and directrix y = 1.

OpenStudy (tw101):

Im not sure how to start.

OpenStudy (tw101):

1?

OpenStudy (tw101):

okay the formula you put up above?

OpenStudy (tw101):

oh okay, onne second and then ill tell you what I got:)

OpenStudy (jdoe0001):

one may note that 1,2 is the focus point, not the vertex point

OpenStudy (tw101):

oh wait what does that mean then?

OpenStudy (jdoe0001):

well... she maybe crazy, but still beautiful, so tis ok =)

OpenStudy (tw101):

haha.

OpenStudy (tw101):

im seriously confused though...

OpenStudy (jdoe0001):



OpenStudy (tw101):

Im not sure... Im sorry, this is new to me.

OpenStudy (jdoe0001):

hmmm well.. have you covered parabolas yet?

OpenStudy (tw101):

Yes, briefly.

OpenStudy (jdoe0001):

well.. it turns out that the focus point is "p" distance from the vertex and the directrix is "p" distance also from the vertex so.. the vertex is really half-way between both, since the focus and directrix are the same distance from the vertex keeping in mind that, the parabola opens up towards the focus point

OpenStudy (tw101):

Okay, well crazyandbeautiful helped me out and said that (1,3/2) was the vertex. haha.

OpenStudy (jdoe0001):

hmmm how did she get 1, 3/2 though?

OpenStudy (tw101):

haha no you dont need to its okay! @Crazyandbeautiful

OpenStudy (tw101):

okay so its always halfway between the focus point and directrix?

OpenStudy (jdoe0001):

yeap, they're always "p" distance from the vertex

OpenStudy (tw101):

Okay, i get that now:)

OpenStudy (tw101):

Where do I go from there?

OpenStudy (jdoe0001):

so... get the vertex, get "p" and plug and chug \(\large{ \begin{array}{llll} (y-{\color{blue}{ k}})^2=4{\color{purple}{ p}}(x-{\color{brown}{ h}}) \\ (x-{\color{brown}{ h}})^2=4{\color{purple}{ p}}(y-{\color{blue}{ k}})\\ \end{array} \qquad \begin{array}{llll} vertex\ ({\color{brown}{ h}},{\color{blue}{ k}})\\ {\color{purple}{ p}}=\textit{distance from vertex to }\\ \qquad \textit{ focus or directrix} \end{array}}\)